An existence result for a semi-positone problem with a sign changing weight

Jaffar Ali Shahul-Hameed Mississippi State University, USA

In this paper we establish an existence result on positive solution for a class of reaction-diffusion equation with semi-positone structure. In particular, our results apply to the diffusive logistic equation with a class of sign changing weight and constant yield harvesting. We establish the result via the method of sub-super solutions.

Joint Work with Dr. R. Shivaji.